# SQL-R1: Training Natural Language to SQL Reasoning Model By Reinforcement Learning

### v4

@@ Abstract
Natural Language to SQL (NL2SQL) enables intuitive interactions with databases by transforming natural language queries into structured SQL statements. Despite recent advancements in enhancing human-computer interaction within database applications, significant challenges persist, particularly regarding the inference performance in complex scenarios involving multi-table joins and nested queries. Current methodologies primarily utilize supervised fine-tuning (SFT) to train the NL2SQL model, which may limit adaptability and interpretability in new environments  (e.g., finance and healthcare). In order to enhance the reasoning performance of the NL2SQL model in the above complex situations, we introduce SQL-R1, a novel NL2SQL reasoning model trained by the reinforcement learning (RL) algorithms. We design a specialized RL-based reward function tailored for NL2SQL tasks and discussed the impact of cold start on the effectiveness of intensive training. In addition, we achieve competitive accuracy using only a tiny amount of synthetic NL2SQL data for augmented training and further explore data engineering for RL. In existing experiments, SQL-R1 achieves execution accuracy of 88.6% and 67.1% on the benchmark Spider and BIRD, respectively.
@@ 1 Introduction
Despite these advancements, improving the NL2SQL system’s inference performance in complex database scenarios remains a considerable challenge. As shown in Figure 1, schema complexity may lead to generation errors in processing multi-table joins and nested queries, and it is difficult for individually trained models to think and process complex semantics. Currently, a significant portion of NL2SQL research is devoted to training open-source large language models (LLMs) by supervised fine-tuning (SFT) [13, 14, 15] to achieve accuracy at a smaller model scale compared to approaches using closed-source LLMs (e.g., GPT-4, GPT-4o) [16, 8, 10]. However, SFT relies on the database schema structure and the training data scale. This may lead to the existing model’s instability in domain adaptation and generalization in new database environments. Additionally, the lack of interpretability of NL2SQL reasoning logic limits the application of the model in high-risk fields, such as finance and healthcare.
@@
Q3Can we deploy sustainable data engineering for training robust and efficient NL2SQL reasoning models? RL training relies on high-quality training data, while current NL2SQL tasks lack a large amount of real data for training. How to develop the data support for NL2SQL inference model based on the existing data engineering technology is an important challenge to solve the model training, improve the robustness and generalization of the model.
@@ 2.4 SQL Candidates Selection
In order to select the most appropriate SQL in the inference process, the model generates several SQL candidates and their thought processes for a problem. We execute all SQL candidates and select the SQL with the highest score as the final answer based on self-consistency voting. Notably, the reasoning response of SQL-R1 comprises an observable process of thinking and interpreting, making the results easier for the user to understand.
@@ 3.2 Main Results
The results presented in Table 1 underscore the performance of SQL-R1 across various foundational models. Specifically, when trained on the Qwen2.5-Coder-3B, SQL-R1 achieved an execution accuracy of 78.1% on the Spider development dataset, 78.9% on the Spider test dataset, and 54.6% on the BIRD development dataset, demonstrating robust performance even with a smaller base model. Moreover, SQL-R1 exhibits a marked enhancement in accuracy when utilized with larger models; for instance, the Qwen2.5-Coder-7B model yielded accuracies of 87.6%, 88.7%, and 63.1% across the same Spider development, Spider test, and BIRD development datasets, respectively. The deployment of the even larger Qwen2.5-Coder-14B model further affirmed SQL-R1’s capabilities, achieving high accuracy rates of 86.7%, 88.1%, and 67.1% on these datasets. In comparison to other solutions leveraging closed-source models, such as GPT-4 and GPT-4o, SQL-R1 demonstrates competitive and often superior performance. Notably, In comparison to contemporaneous exploration of reasoning-based NL2SQL (e.g., DeepRetrieval [31] and Reasoning-SQL [32]), SQL-R1 surpasses them on the Spider-Dev and BIRD-Dev datasets. This performance highlights the efficacy of our proposed approach in utilizing diverse base models to attain state-of-the-art results on complex NL2SQL reasoning tasks. Additional experimental results on other NL2SQL benchmarks (e.g., Spider2.0) can be found in Table 4, Table 5 and Section B.1.
@@
3.3 Analysis of SFT Cold Start
@@
In addressing Q3, experimental results corroborate the assertion that synthetic data engineering plays a pivotal role in augmenting the inference capabilities of NL2SQL models, especially when utilized in conjunction with RL. The synergistic effect of synthetic data and RL training reveals marked enhancements in model performance. For instance, SQL-R1, when trained on SynSQL-Complex-5K, surpasses the performance of both the 7B and 14B baseline models. This finding elucidates the substantial potential inherent in synthetic data engineering to enhance model capabilities. Ongoing research endeavors will persist in exploring and refining strategies for the integration of synthetic data into the training regimen, with the aim of further advancing model performance and generalization.
3.4 Ablation Study of Reward Components
We conducted an ablation experiment on the BIRD development dataset to verify the proposed reinforcement learning reward mechanism’s effectiveness. The experiment sequentially removed individual reward components from the comprehensive reward function while maintaining the parameter settings established in Section 3.1. As presented in Table 4, the result demonstrates that omitting any reward component from the original reward function adversely impacts inference performance. This underscores the critical importance of execution feedback and result reward in the model training process. Section B.4 and B.5 provide more details of the ablation study.
Figure 2: Performance and model scale on the BIRD-Dev dataset.
@@ 4 Related Works
Recent research on reinforcement learning for LLMs has increasingly concentrated on enhancing reasoning capabilities and optimizing interaction with the external environment to improve proficiency in complex, multi-step reasoning tasks [48, 49]. In recent months, DeepSeek-R1 [17] have illustrated the effectiveness of group relative policy optimization (GRPO), a novel and efficient reinforcement algorithm in cultivating robust reasoning behaviors in LLMs. By focusing exclusively on the correctness of the final answer, GRPO facilitates the internalization of intermediate reasoning steps without the necessity of explicit supervision, achieving notable success across various domains, including mathematics [50], finance [18], programming [51], user interface [52, 53] and vision tasks [54]. Concurrently, some studies have explored integrating external tools, such as retrieval [31] and search engines [19], into reinforcement learning frameworks for LLM reasoning, primarily focusing on training models to effectively employ a single, general-purpose tool to support specific reasoning chains. In light of these advancements, we aim to extend LLM reasoning capabilities to the NL2SQL task by adopting a reinforcement learning approach that dynamically adjusts the model’s decision-making strategy through interaction with the environment, thereby enhancing its ability to manage complex semantics and improve generalization and domain adaptation capabilities.
@@ Appendix B Additional Results and Analysis
NL2SQL Method Base Model Spider-DK Spider-Syn Spider-Realistic
SENSE [55] CodeLlama-7B 77.9 72.6 82.7
ROUTE [56] Llama3-8B 74.6 77.4 80.9
SQL-o1 [30] Llama3-8B 78.7 72.6 82.7
OmniSQL [22] Qwen2.5-Coder-7B 77.8 69.6 78.0
SQL-PaLM [57] PaLM-2 67.5 70.9 77.4
PURPLE [58] GPT-4 75.3 74.0 79.9
SQL-R1 (Ours) Qwen2.5-Coder-3B 70.5 66.4 71.5
SQL-R1 (Ours) Qwen2.5-Coder-7B 78.1 76.7 83.3
SQL-R1 (Ours) Qwen2.5-Coder-14B 79.3 78.5 86.2
NL2SQL Method Base Model Simple Moderate Challenging All
CodeS Codes-15B 65.8 48.8 42.4 58.5
DAIL-SQL GPT-4 63.0 45.6 43.1 55.9
SuperSQL GPT-4 66.9 46.5 43.8 58.5
SQL-R1 (Ours) Qwen2.5-Coder-7B 72.1 60.8 51.0 66.6
SQL-R1 (Ours) Qwen2.5-Coder-14B 72.4 59.7 56.5 67.1
B.1 Additional Analysis of Performance on NL2SQL Benchmarks
Performance on other NL2SQL Benchmarks.
Table 5 demonstrates the comparison of execution accuracy on Spider-DK, Spider-Syn and Spider-Realistic dataset. Notably, SQL-R1 exhibits superior performance relative to above NL2SQL solutions, which further prove the generalization and robustness of a under different database and query requirements.
Complexity Analysis.
B.2 Analysis and Comparison of Base LLMs
@@
B.3 Analysis and Comparison of Efficiency
@@
B.4 Ablation Study of SQL Candidate Rollouts in Reasoning
To assess the efficacy of the self-consistency in the inference, we conducted ablation studies on the Qwen2.5-Coder-3B and Qwen2.5-Coder-7B base models, aiming to determine the optimal number of SQL candidate rollouts generated under varied temperature settings designed to enhance diversity. The results illustrated in Figure 9 reveal that setting the number of SQL candidates to 8 strikes the ideal balance between diversity and computational efficiency. In the case of the 3B model, configurations with fewer than eight candidates resulted in diminished execution accuracy due to insufficient diversity, whereas exceeding 8 candidates led to a plateau in accuracy alongside increased computational overhead. The 7B model demonstrated relatively stable accuracy across different candidate counts but achieved its highest accuracy at the same optimal setting. These findings substantiate that the parameter configuration of 8 SQL candidates and a temperature setting of 0.8 effectively enhances SQL-R1’s performance, successfully navigating the trade-off between candidate diversity and computational cost, thereby reinforcing the model’s robustness in SQL query generatio
@@ B.4 Ablation Study of SQL Candidate Rollouts in Reasoning
B.5 Ablation Study of Database Value Retrieval for RL Training
@@
B.6 Ablation Study of Reward Sensitivity
@@
B.7 Visualization of Training Process

### v5

@@ Abstract
Natural Language to SQL (NL2SQL) enables intuitive interactions with databases by transforming natural language queries into structured SQL statements. Despite recent advancements in enhancing human-computer interaction within database applications, significant challenges persist, particularly regarding the reasoning performance in complex scenarios involving multi-table joins and nested queries. Current methodologies primarily utilize supervised fine-tuning (SFT) to train the NL2SQL model, which may limit adaptability and interpretability in new environments  (e.g., finance and healthcare). In order to enhance the reasoning performance of the NL2SQL model in the above complex situations, we introduce SQL-R1, a novel NL2SQL reasoning model trained by the reinforcement learning (RL) algorithms. We design a specialized RL-based reward function tailored for NL2SQL tasks and discussed the impact of cold start and synthetic data on the effectiveness of intensive training. In addition, we achieve competitive accuracy using only a tiny amount of synthetic NL2SQL data for augmented training and further explore data engineering for RL. In existing experiments, SQL-R1 achieves execution accuracy of 88.6% and 67.1% on the benchmark Spider and BIRD, respectively. The code is available at https://github.com/IDEA-FinAI/SQL-R1.
@@ 1 Introduction
Despite these advancements, improving the NL2SQL system’s reasoning performance in complex database scenarios remains a considerable challenge. As shown in Figure 1, schema complexity may lead to generation errors in processing multi-table joins and nested queries, and it is difficult for individually trained models to think and process complex semantics. Currently, a significant portion of NL2SQL research is devoted to training open-source large language models (LLMs) by supervised fine-tuning (SFT) [13, 14, 15] to achieve accuracy at a smaller model scale compared to approaches using closed-source LLMs (e.g., GPT-4, GPT-4o) [16, 8, 10]. However, SFT relies on the database schema structure and the training data scale. This may lead to the existing model’s instability in domain adaptation and generalization in new database environments. Additionally, the lack of interpretability of NL2SQL reasoning logic limits the application of the model in high-risk fields, such as finance and healthcare.
@@
Q3Can we deploy sustainable data engineering for training robust and efficient NL2SQL reasoning models? RL training relies on high-quality training data, while current NL2SQL tasks lack a large amount of real data for training. How to develop the data support for NL2SQL reasoning model based on the existing data engineering technology is an important challenge to solve the model training, improve the robustness and generalization of the model.
@@ 2.4 SQL Candidates Selection
In order to select the most appropriate SQL in the reasoning process, the model generates several SQL candidates and their thought processes for a problem. We execute all SQL candidates and select the SQL with the highest score as the final answer based on self-consistency voting. Notably, the reasoning response of SQL-R1 comprises an observable process of thinking and interpreting, making the results easier for the user to understand.
@@ 3.2 Main Results
The results presented in Table 1 underscore the performance of SQL-R1 across various foundational models. Specifically, when trained on the Qwen2.5-Coder-3B, SQL-R1 achieved an execution accuracy of 78.1% on the Spider development dataset, 78.9% on the Spider test dataset, and 54.6% on the BIRD development dataset, demonstrating robust performance even with a smaller base model. Moreover, SQL-R1 exhibits a marked enhancement in accuracy when utilized with larger models; for instance, the Qwen2.5-Coder-7B model yielded accuracies of 87.6%, 88.7%, and 63.1% across the same Spider development, Spider test, and BIRD development datasets, respectively. The deployment of the even larger Qwen2.5-Coder-14B model further affirmed SQL-R1’s capabilities, achieving high accuracy rates of 86.7%, 88.1%, and 67.1% on these datasets. In comparison to other solutions leveraging closed-source models, such as GPT-4 and GPT-4o, SQL-R1 demonstrates competitive and often superior performance. Notably, In comparison to contemporaneous exploration of reasoning-based NL2SQL (e.g., DeepRetrieval [31] and Reasoning-SQL [32]), SQL-R1 surpasses them on the Spider-Dev and BIRD-Dev datasets. This performance highlights the efficacy of our proposed approach in utilizing diverse base models to attain state-of-the-art results on complex NL2SQL reasoning tasks.
Table 2 demonstrates the comparison of execution accuracy on Spider-DK, Spider-Syn and Spider-Realistic dataset. Notably, SQL-R1 exhibits superior performance relative to above NL2SQL solutions, which further prove the generalization and robustness of a under different database and query requirements. Additional experimental results on other NL2SQL benchmarks (e.g., Spider2.0) can be found in Table 6 and Section B.
NL2SQL Method Base Model Spider-DK Spider-Syn Spider-Realistic
SENSE [37] CodeLlama-7B 77.9 72.6 82.7
ROUTE [38] Llama3-8B 74.6 77.4 80.9
SQL-o1 [30] Llama3-8B 78.7 72.6 82.7
OmniSQL [22] Qwen2.5-Coder-7B 77.8 69.6 78.0
SQL-PaLM [39] PaLM-2 67.5 70.9 77.4
PURPLE [40] GPT-4 75.3 74.0 79.9
SQL-R1 (Ours) Qwen2.5-Coder-3B 70.5 66.4 71.5
SQL-R1 (Ours) Qwen2.5-Coder-7B 78.1 76.7 83.3
SQL-R1 (Ours) Qwen2.5-Coder-14B 79.3 78.5 86.2
NL2SQL Method Base Model Simple Moderate Challenging All
CodeS [14] Codes-15B 65.8 48.8 42.4 58.5
DAIL-SQL [16] GPT-4 63.0 45.6 43.1 55.9
SuperSQL [33] GPT-4 66.9 46.5 43.8 58.5
SQL-R1 (Ours) Qwen2.5-Coder-7B 72.1 60.8 51.0 66.6
SQL-R1 (Ours) Qwen2.5-Coder-14B 72.4 59.7 56.5 67.1
Complexity Analysis.
@@
Figure 2: Performance and model scale on the BIRD-Dev dataset.
@@ 3.3 Analysis of SFT Cold Start
In addressing Q3, experimental results corroborate the assertion that synthetic data engineering plays a pivotal role in augmenting the reasoning capabilities of NL2SQL models, especially when utilized in conjunction with RL. The synergistic effect of synthetic data and RL training reveals marked enhancements in model performance. For instance, SQL-R1, when trained on SynSQL-Complex-5K, surpasses the performance of both the 7B and 14B baseline models. This finding elucidates the substantial potential inherent in synthetic data engineering to enhance model capabilities. Ongoing research endeavors will persist in exploring and refining strategies for the integration of synthetic data into the training regimen, with the aim of further advancing model performance and generalization.
@@
3.4 Ablation Study of Reward Components
We conducted an ablation experiment on the BIRD development dataset to verify the proposed reinforcement learning reward mechanism’s effectiveness. The experiment sequentially removed individual reward components from the comprehensive reward function while maintaining the parameter settings established in Section 3.1. As presented in Table 6, the result demonstrates that omitting any reward component from the original reward function adversely impacts reasoning performance. This underscores the critical importance of execution feedback and result reward in the model training process. Section B.3 and B.4 provide more details of the ablation study.
@@ 4 Related Works
Recent research on reinforcement learning for LLMs has increasingly concentrated on enhancing reasoning capabilities and optimizing interaction with the external environment to improve proficiency in complex, multi-step reasoning tasks [52, 53, 54]. In recent months, DeepSeek-R1 [17] have illustrated the effectiveness of group relative policy optimization (GRPO), a novel and efficient reinforcement algorithm in cultivating robust reasoning behaviors in LLMs. By focusing exclusively on the correctness of the final answer, GRPO facilitates the internalization of intermediate reasoning steps without the necessity of explicit supervision, achieving notable success across various domains, including mathematics [55], finance [18], programming [56], user interface [57, 58] and vision tasks [59]. Concurrently, some studies have explored integrating external tools, such as retrieval [31] and search engines [19], into reinforcement learning frameworks for LLM reasoning, primarily focusing on training models to effectively employ a single, general-purpose tool to support specific reasoning chains. In light of these advancements, we aim to extend LLM reasoning capabilities to the NL2SQL task by adopting a reinforcement learning approach that dynamically adjusts the model’s decision-making strategy through interaction with the environment, thereby enhancing its ability to manage complex semantics and improve generalization and domain adaptation capabilities.
@@ Appendix B Additional Results and Analysis
B.1 Analysis and Comparison of Base LLMs
@@
B.2 Analysis and Comparison of Efficiency
@@
B.3 Ablation Study of SQL Candidate Rollouts in Reasoning
To assess the efficacy of the self-consistency in the inference, we conducted ablation studies on the Qwen2.5-Coder-3B and Qwen2.5-Coder-7B base models, aiming to determine the optimal number of SQL candidate rollouts generated under varied temperature settings designed to enhance diversity. The results illustrated in Figure 9 reveal that setting the number of SQL candidates to 8 strikes the ideal balance between diversity and computational efficiency. In the case of the 3B model, configurations with fewer than eight candidates resulted in diminished execution accuracy due to insufficient diversity, whereas exceeding 8 candidates led to a plateau in accuracy alongside increased computational overhead. The 7B model demonstrated relatively stable accuracy across different candidate counts but achieved its highest accuracy at the same optimal setting. These findings substantiate that the parameter configuration of 8 SQL candidates and a temperature setting of 0.8 effectively enhances SQL-R1’s performance, successfully navigating the trade-off between candidate diversity and computational cost, thereby reinforcing the model’s robustness in SQL query generation.
@@ B.3 Ablation Study of SQL Candidate Rollouts in Reasoning
B.4 Ablation Study of Database Value Retrieval for RL Training
@@
B.5 Ablation Study of Reward Sensitivity
@@
B.6 Visualization of Training Process